OpenStudy (anonymous):

(Help please) Gregor is documenting the height of pea plants each week. He has determined the function to be f(x) = 2x + 1. Based on the situation above, describe any necessary restrictions to the domain and range. Provide justification for the restrictions.

OpenStudy (anonymous):

if it is weeks, since we cannot go backwards in time, i would say the domain is \(x\geq 0\)

OpenStudy (anonymous):

Oh that makes sense, thank you.

OpenStudy (anonymous):

you might want to put a cap on it, since presumably the pea plants do not live forever, maybe 2 years? maybe only one? i suppose you could say \[0\leq x \leq 52\] if they are annuals, not sure how long they live

OpenStudy (anonymous):

as for the range, since the begging is \(x=0\) you have \(y=1\) so maybe \(y\geq 1\) for the range, or put a cap on that as well, since they cannot get infinitely tall

OpenStudy (anonymous):

*beginning

OpenStudy (anonymous):

ok, Thanks so much, that was really helpful.

OpenStudy (anonymous):

yw